OpenStudy (anonymous):

how do i make 0.1515.... in to a fraction

OpenStudy (anonymous):

First, count how many are not repeating: 0

OpenStudy (anonymous):

Subtract the number from the whole decimal

OpenStudy (anonymous):

x = 0.1515... 100x = 15.1515... 100x - x = 15

OpenStudy (anonymous):

99x = 15 x = 15/99 x = 5/33